# Going beyond the creative limits of next-token prediction

### v3

@@ 1 Introduction
Not all forms of intelligence are solely about being correct or wrong. In open-ended tasks, what also matters is finding creative ways to satisfy a request, making surprising and fresh connections never seen before. For instance, consider responding to highly under-specified prompts like “Generate a challenging high-school word problem involving the Pythagoras Theorem.” or “Suggest some candidate therapeutic antibodies targeting the HER2 antigen.” or “Provide a vivid analogy to differentiate quantum and classical mechanics.” Creativity in such tasks requires generating responses that are not just correct or coherent, but are also diverse across responses and are original compared to the training data. These currently-sidelined desiderata will rise to prominence as we explore LLMs for open-ended scientific discovery (Gruver et al., 2023; Romera-Paredes et al., 2024; Si et al., 2024; Lu et al., 2024a), for generating novel training data (Yu et al., 2024; Yang et al., 2024c; Wang et al., 2023), and as we scale up test-time compute approaches that benefit from diversity in exploration, such as best-of-N (Cobbe et al., 2021; Chow et al., 2024; Dang et al., 2025) and long chain-of-thought reasoning (OpenAI, 2024; DeepSeek-AI, 2025; Snell et al., 2024; Wu et al., 2024).
Unlike simple open-ended tasks like generating names and basic sentences (Zhang et al., 2024b; Hopkins et al., 2023), many creative tasks (like designing a clever Olympiad problem) are said to involve a random flash of creative insight termed variously as a leap of thought (Wang et al., 2024a; Talmor et al., 2020; Zhong et al., 2024), a “eureka” moment (Bubeck et al., 2023), a mental leap (Holyoak & Thagard, 1995; Callaway, 2013; Hofstadter, 1995) or an incubation step (Varshney et al., 2019). The thesis of this paper is that learning to solve such creative leap-of-thought tasks (defined shortly) is misaligned with the current language modeling paradigm (a) in terms of next-token learning, and (b) in how randomness is elicited. We articulate these two concerns by designing a suite of algorithmic tasks inspired by such creative tasks. We then demonstrate how the creativity of language models suffers in these tasks, and how this can be alleviated (to an extent, within our tasks).
@@
In search of more definitive conclusions, we approach from a different angle: we study minimal and controllable tasks that are loose abstractions of real-world tasks and yet allow one to rigorously quantify originality and diversity. This follows along the lines of recent works that have studied the diversity of models in graph path-finding (Khona et al., 2024) and generating challenging context-free grammars (CFGs) (Allen-Zhu & Li, 2023b). Broadly, we refer to such tasks as open-ended algorithmic tasks. Our aim is to design tasks more minimal than these prior tasks, and crucially, tease apart distinct computational skills required for creativity. This will allow us to systematically investigate issues in the current paradigm of model training and propose alternatives.
@@
Next, we turn to how we elicit randomness from a Transformer. While the de facto approach is to elicit randomness at the output — temperature sampling — we contrast this with injecting randomness into the model. Concretely, we study seed-conditioning, where we train with a random prefix string per example. Surprisingly, we find that seed-conditioning induces non-trivial algorithmic creativity even with deterministic, greedy decoding. In fact, seed-conditioning is comparable with and in some cases, more creative than temperature sampling. Intuitively, maximizing diversity at the output-token-level is computationally burdensome: it requires simultaneously processing a diverse set of leaps of thoughts to compute a marginalized token distribution. It is easier to first sample a single latent leap of thought, and then compute the token conditioned on that one leap. We conjecture that seed-conditioning enables a latent-leap-conditioned token generation.
Overall, we hope our study advances the field in two directions. First, we provide a new angle to advocate for multi-token approaches, orthogonal to the “path-star” example in Bachmann & Nagarajan (2024) (or B&N’24 in short). Whereas, the path-star example portrays a gap in correctness of reasoning, ours shows a gap in diversity of open-ended thinking. We note though that B&N’24 is an impossibility result where next-token learning breaks down spectacularly (unless there is exponential data or compute), while ours is a data-inefficiency result (where next-token learning occurs but is mediocre). Next, the gap we show appears even in 2-token-lookahead tasks as against the many-token-lookahead path-star task. Third, and perhaps most conceptually important is the fact that, while the path-star task is amenable to next-token prediction upon reversing the tokens, we identify tasks where no re-ordering is friendly towards next-token prediction — the optimal thing to do is to globally learn higher-order patterns implicit in the whole future sequence. This presents a challenge to recent proposals that aim to fix next-token prediction via permutations (Pannatier et al., 2024; Thankaraj et al., 2025) or partial lookaheads (Bavarian et al., 2022; Fried et al., 2022; Kitouni et al., 2024; Nolte et al., 2024).
@@
    We find that multi-token prediction through teacherless training or diffusion, results in significantly increased algorithmic creativity in our tasks compared to next-token prediction.
@@ 2 Open-ended algorithmic tasks & two types of creativity
Open-ended tasks that do not require a leap of thought. One simple open-ended task that may come to mind is generating uniformly-random known entities, like celebrity names (Zhang et al., 2024b). However, there is no opportunity to create a novel string here. A more interesting example may be generating grammatically coherent probabilistic CFG strings following a subject verb object format e.g., the cat chased a rat (Hopkins et al., 2023). While novel strings become possible now, no sophisticated leaps of thought are involved; each token can be generated on the fly, satisfying a local next-token constraint to be coherent.
In light of this, we can rephrase our goal as designing open-ended, creative tasks where coherence requires satisfying more interesting, “global” constraints. To build this systematically, we draw inspiration from literature in cognitive science (Boden, 2003). Boden (2003) argues that fundamentally, there are three forms of creativity in that order: combinational, exploratory and transformative. We elaborate on the first two (the last, we do not look at).
@@ 2.1 The fundamental types of creativity (Boden, 2003)
In the upcoming sections, we will attempt to capture some core computations of rudimentary instances within the two classes of creative skills above. We emphasize that by no means does our minimal algorithmic setup intend to capture the human values or the subjective aspects of these endeavors (Schmidhuber, 2009); nor do they capture the rich array of creative acts that Boden (2003) discusses within these categories. (See limitations in §6).
@@ 2.2 The basic setting and notations
In all our tasks, we assume the standard generative model setting: the model must learn an underlying distribution $\mathcal{D}$ through a training set $S$ of $m$ independent samples $\bm{s}_{i}\sim\mathcal{D}$. The distribution is over a space $\mathbb{V}^{L}$ of $L$-length strings composed of tokens from the vocabulary $\mathbb{V}$. The tasks are open-ended in that there is no one correct answer at test-time. The goal is to produce a random string from $\mathcal{D}$, much like responding to the query Design a high-school word problem. We note that this setup resembles how models are trained in pretraining or in applications like drug discovery or protein modeling.
@@
Algorithmic creativity: Upon witnessing a finite set of examples, the model must learn to generate only strings that are (a) coherent, (b) original (not memorized) and (c) diverse (covers the whole support). An exact quantification of this is computationally expensive in our tasks. Instead, we approximate it by sampling a set $T$ of many independent generations from the model and computing the fraction of $T$ that is original, coherent and unique.
Let the boolean $\texttt{mem}_{S}(\bm{s})$ denote whether an example $\bm{s}$ is from the training set $S$ and let the integer function $\texttt{uniq}(X)$ denote the number of unique examples in a set $X$. (The exact definitions of these quantities vary by tasks, as we will see). Then, we define our (empirical) algorithmic creativity metric:
@@
Our setup models an in-distribution form of novelty as it offers a rigorous way to study the problem. Admittedly, this is a far simpler form of novelty than what is expected in the real-world. Nevertheless, even this simple setting will help us appreciate the limits of current language models.
@@ 2.3.1 Sibling discovery
A note on the leap of thought. To concretely illustrate what we mean by a leap of thought, we note that the above task can be designed with or without a leap. Observe that the most natural order of generation is to generate the parent vertex (i.e., punchline) first, and pick the siblings after (conditioned on the parent). Thus, if the task demanded the ordering $(\Gamma,\gamma,\gamma^{\prime})$, it would involve no leap of thought: each next token can be learned and generated through simple rules conditioned on the past, without planning.
However, the word play example involves a non-sequential leap of thought in that even though the punchline (the parent) appears last, it must be planned ahead of time. Paralleling this leap-of-thought structure, we define our sibling discovery task to generate the triplets as $\bm{s}=(\gamma,\gamma^{\prime},\Gamma$), where the siblings appear first. We hypothesize that this (sibling-first) construction is adversarial towards next-token learning, while a reversed (parent-first) dataset is friendlier towards next-token learning. More on this in §2.6.
@@ 2.4.2 Line construction
2.5 Permutation-invariance of our tasks
We emphasize a key novel aspect in our last three tasks. Many algorithmic tasks in literature like addition (Lee et al., 2024), or the path-star task (B&N’24) or Sibling Discovery have a natural ordering in which the tokens can be learned and generated, even if not left to right. However, the Triangle Discovery, Line Construction and Circle Construction tasks are permutation-invariant — no token is more privileged than the other, and all tokens must be “simultaneously learned” to infer the underlying process. We view this as an abstraction of real-world tasks where the creative process is highly implicit, and not immediate from the text. These tasks offer a test-bed even for non-next-token approaches that rely on re-permuting the tokens (Pannatier et al., 2024; Thankaraj et al., 2025) or predicting only a part of the future (Kitouni et al., 2024; Nolte et al., 2024; Bavarian et al., 2022; Fried et al., 2022).
@@ 2.6 How next-token learning may suffer in our tasks
Much like in sophisticated creative tasks, in our tasks, the most natural way to generate the string is by planning various random latent choices (say $\bm{z}$) in advance and by producing a plan-conditioned distribution $p(\bm{s}|\bm{z})$ over coherent strings $\bm{s}$. However, next-token prediction (NTP) – or next-token learning to be precise – we argue, is myopic and fails to learn such a latent plan. Our argument extends that of B&N’24 to our even simpler tasks.
@@
Consider learning the Sibling Discovery where we must generate sibling-parent triplets $(\gamma,\gamma^{\prime},\Gamma)$. Even if the parent must be emitted last, the most natural generative rule is to plan the parent first and decide the children last. We can think of this as learning a latent plan $\bm{z}:=\Gamma$. Then, learning the plan-conditioned generation $p(\gamma,\gamma^{\prime},\Gamma|\bm{z})$ factorizes to learning the distribution of children conditioned on a parent as $p(\gamma|\bm{z}:=\Gamma)$ and $p(\gamma^{\prime}|\bm{z}:=\Gamma)$ (due to conditional independence), and the trivial $p(\Gamma|\bm{z}:=\Gamma)$. This requires only as many parent-sibling edges as there are in the graph, i.e., $O(m\cdot n)$ many points, if there are $m$ parents, each with $n$ children. This is optimal.
@@
Things proceed differently with NTP. We argue that a NTP-learner would fail to learn the plan $\bm{z}:=\Gamma$. The key intuition is that that an NTP-learner learns the parent $\Gamma$ witnessing the siblings $(\gamma,\gamma^{\prime})$ as input. This is trivial to fit: the parent is simply the mutual neighbor of the two siblings revealed in the prefix! B&N’24 term such shortcuts as Clever Hans cheats since the model witnesses and exploits part of the ground-truth it must generate (the siblings). Such cheats are simpler than even the true generative rule and are thus quickly picked up during learning. The model then loses any supervision to learn the latent plan, $z:=\Gamma$.
@@
After the Clever Hans cheat is learned, the NTP-learner learns the second sibling not through the plan-conditioned distribution $p(\gamma^{\prime}|\bm{z}:=\Gamma)$ but through the next-token-conditional, $p(\gamma^{\prime}|\gamma)$. This is a complex distribution: learning this would require witnessing every sibling-sibling pair totalling $O(m\cdot n^{2})$ many training data — larger by a factor of $n$ than the data requirements of the more natural rule.
@@
More abstractly, in our tasks, it is most efficient to learn a well-planned random latent $p(\bm{z})$ and a subsequent latent-conditioned distribution $p(\bm{s}|\bm{z})$. However, NTP factorizes this into pieces of the form $p(s_{i}|\bm{s}_{<i},\bm{z})$. Consequently, the model learns uninformative latents from the later tokens, lured by Clever Hans cheats. Conversely, the earlier tokens are learned through complex rules bereft of a latent plan. While this may not lead to complete breakdown of learning as in B&N’24, it must lead to data-hungry learning.

### v4

@@ 1 Introduction
Not all forms of intelligence are solely about being correct or wrong. In open-ended tasks, what also matters is finding creative ways to satisfy a request, making surprising and fresh connections never seen before. For instance, consider responding to highly under-specified prompts like “Generate a challenging high-school word problem involving the Pythagoras Theorem.” or “Suggest some candidate therapeutic antibodies targeting the HER2 antigen.” or “Provide a vivid analogy to differentiate quantum and classical mechanics.” Creativity in such tasks requires generating responses that are not just correct or coherent, but are also diverse across responses and are original compared to the training data. These currently-sidelined desiderata will rise to prominence as we explore LLMs for open-ended scientific discovery (Gruver et al., 2023; Romera-Paredes et al., 2024; Hayes et al., 2025), for generating novel training data (Yu et al., 2024; Yang et al., 2024c; Wang et al., 2023), and as we scale up test-time compute approaches that benefit from diversity in exploration, such as best-of-N (Cobbe et al., 2021; Chow et al., 2024; Dang et al., 2025) and long chain-of-thought reasoning (OpenAI, 2024; DeepSeek-AI, 2025; Snell et al., 2024; Wu et al., 2024).
Unlike simple open-ended tasks like generating names and basic sentences (Zhang et al., 2024b; Hopkins et al., 2023; Bigelow et al., 2024), many creative tasks (like designing a clever Olympiad problem) are said to involve a random flash of creative insight termed variously as a leap of thought (Wang et al., 2024a; Talmor et al., 2020; Zhong et al., 2024), a “eureka” moment (Bubeck et al., 2023), a mental leap (Holyoak & Thagard, 1995; Callaway, 2013; Hofstadter, 1995) or an incubation step (Varshney et al., 2019). The thesis of this paper is that learning to solve such creative leap-of-thought tasks (defined shortly) is misaligned with the current language modeling paradigm (a) in terms of next-token learning, and (b) in how randomness is elicited. We articulate these two concerns by designing a suite of algorithmic tasks inspired by such creative tasks. We then demonstrate how the creativity of language models suffers in these tasks, and how this can be alleviated (to an extent, within our tasks).
@@
In search of more definitive conclusions, we approach from a different angle: we study controllable tasks that are loose abstractions of real-world tasks and yet allow one to objectively quantify originality and diversity. This follows recent works that have studied diversity of models in path-finding (Khona et al., 2024) and challenging context-free grammars (CFGs) (Allen-Zhu & Li, 2023b). We broadly term these as open-ended algorithmic tasks. Our aim then is design minimal instances of these tasks, akin to basic arithmetic for reasoning, so as to tease apart the bare minimum computational skills required for creativity. Then, we can examine how the current modeling paradigm may be limited even with this basic skills, and what alternative paradigms may be desired.
@@
Next, we turn to how we elicit randomness from a Transformer. While the de facto approach is to elicit randomness at the output — temperature sampling — we contrast this with injecting randomness into the model. Concretely, we study seed-conditioning, where we train and test with a random prefix string per example. Surprisingly, we find that seed-conditioning induces non-trivial algorithmic creativity even with deterministic decoding; it is in fact, comparable to temperature sampling (and in some cases, better). We intuit that maximizing diversity at the output-level results in cognitive overload: it requires simultaneously processing many leaps of thoughts to compute a marginalized token distribution. Input-level noise injection could sidestep this by articulating one leap of thought per seed. We also view this exploration as a controlled study solidifying prior indications that prompt variations induce diversity (Li et al., 2023; Lau et al., 2024; Naik et al., 2024; Li et al., 2022).
Overall, we hope to advance the field in two directions. First, we provide a new angle to advocate for multi-token approaches, orthogonal to the “path-star” example in Bachmann & Nagarajan (2024) (or B&N’24 in short). Whereas, the path-star example portrays a gap in correctness of reasoning, ours shows a gap in diversity of open-ended thinking. We note though that B&N’24 is an impossibility result where next-token learning breaks down spectacularly (unless there is exponential data or compute), while ours is a data-inefficiency result (where next-token learning occurs but is mediocre). Next, the gap we show appears even in 2-token-lookahead tasks as against the many-token-lookahead path-star task. Third, and most conceptually important is the fact that, while the path-star task is amenable to next-token prediction upon reversing the tokens, we identify tasks where no re-ordering is friendly towards next-token prediction — the optimal thing to do is to globally learn higher-order patterns implicit in the whole future sequence. This presents a challenge to recent proposals that aim to fix next-token prediction via permutations (Pannatier et al., 2024; Thankaraj et al., 2025) or partial lookaheads (Bavarian et al., 2022; Fried et al., 2022; Kitouni et al., 2024; Nolte et al., 2024).
@@
    We find that multi-token prediction through one of teacherless training or diffusion, results in significantly increased algorithmic creativity in our tasks compared to next-token prediction.
@@ 2 Open-ended algorithmic tasks & two types of creativity
Open-ended tasks that do not require a leap of thought. One simple open-ended task that may come to mind is generating uniformly-random known entities, like celebrity names (Zhang et al., 2024b). However, there is no opportunity to create a novel string here. A more interesting example may be generating grammatically coherent probabilistic CFG strings following a subject verb object format e.g., the cat chased a rat (Hopkins et al., 2023). While novel strings become possible now, no sophisticated leaps of thought are involved; each token can be generated on the fly, satisfying a local next-token constraint to be coherent. In light of this, we can rephrase our goal as designing open-ended, creative tasks where coherence requires satisfying more interesting, “global” constraints. To build this systematically, we draw inspiration from literature in cognitive science (Boden, 2003). Boden (2003) argues that fundamentally, there are three forms of creativity in that order: combinational, exploratory and transformative. We elaborate on the first two (the last, we do not look at).
@@ 2.1 The fundamental types of creativity (Boden, 2003)
In the upcoming sections, we will capture some core computations of rudimentary instances within the two creative skills above. By no means does this minimal algorithmic setup intend to capture the human values or the subjective aspects of these endeavors (Schmidhuber, 2009); nor do they capture the rich array of creative acts that Boden (2003) discusses within these categories. (See limitations in §6).
@@ 2.2 The basic setting and notations
In all our tasks, we assume the standard generative model setting: the model must learn an underlying distribution $\mathcal{D}$ through a training set $S$ of $m$ independent samples $\bm{s}_{i}\sim\mathcal{D}$. The distribution is over a space $\mathbb{V}^{L}$ of $L$-length strings composed of tokens from the vocabulary $\mathbb{V}$. The tasks are open-ended in that there is no one correct answer at test-time. The goal is to produce a random string from $\mathcal{D}$, much like responding to the query Design a high-school word problem. We hope that this setup loosely resembles pretraining for language or in applications like drug discovery or protein (Meier et al., 2021; Madani et al., 2023; Watson et al., 2023) and genome modeling (Rives et al., 2021; Nguyen et al., 2024), where the model is trained on a set of varied example sequences, and is later used to produce novel outputs during inference.
@@
Algorithmic creativity: Upon witnessing a finite set of examples, the model must learn to generate only strings that are (a) coherent, (b) original (not memorized) and (c) diverse (covers the whole support). An exact quantification of this is computationally expensive in our tasks. Instead, we approximate it by sampling a set $T$ of many independent generations from the model and computing the fraction of $T$ that is original, coherent and unique. Let the boolean $\texttt{mem}_{S}(\bm{s})$ denote whether an example $\bm{s}$ is from the training set $S$ and let the integer function $\texttt{uniq}(X)$ denote the number of unique examples in a set $X$. (The exact definitions of these quantities vary by tasks, as we will see). Then, we define our (empirical) algorithmic creativity metric:
@@
Admittedly, we are looking at a simple form of novelty that is in-distribution. This notion, while simple, finds its relevance in applications like modeling proteins and genomes. We will also see that this notion is non-trivial enough to demarcate the limits of language models.
@@ 2.3.1 Sibling discovery
What is a leap of thought? The natural order of generation is to pick the parent vertex (i.e., punchline) first, and the siblings after (conditioned on the parent). Thus, if the task demanded producing $(\Gamma,\gamma,\gamma^{\prime})$, the tokens we observe betray the way they are naturally generated. However, the wordplay example presents the punchline (the parent, $\Gamma$) last while implicitly it must be planned ahead. We term this implicit (random) planning step a leap of thought. Paralleling this, we define our sibling discovery task to generate the triplets as $\bm{s}=(\gamma,\gamma^{\prime},\Gamma$), where the parent appears last. We hypothesize that this (sibling-first) construction is adversarial towards next-token learning, while a reversed (parent-first) dataset is friendlier towards next-token learning. (More on this in §2.6.)
@@ 2.4.2 Line construction
2.5 Leap of thought is obscured at the token-level
Rarely does human-written creative output come annotated with an articulation of the background creative process. Nor does each protein or molecule come with an enumeration of the laws that generated it — in fact, it is because those laws are not fully known do we delegate the task of inferring those to a machine. Likewise, in our setting, the training signal consists only of the observed tokens; the model must infer the underlying leap of thought.
In fact, in our last three tasks, the model must infer a deeper form of implicit structure than what exists in many typical algorithmic tasks, like addition (Lee et al., 2024), or the path-star task (B&N’24) or Sibling Discovery. Whereas in those tasks, the leap of thought is perceptible at the token-level, modulo some re-ordering, our last three tasks have no such token ordering. These task are permutation-invariant — no token is more privileged than the other, and all tokens must be “simultaneously learned” to infer the underlying process. We view this as an abstraction of tasks where the creative process is not immediate from the surface of the text. These tasks offer a test-bed even for non-next-token approaches that rely on re-permuting the tokens (Pannatier et al., 2024; Thankaraj et al., 2025) or predicting only a part of the future (Kitouni et al., 2024; Nolte et al., 2024; Bavarian et al., 2022; Fried et al., 2022).
@@ 2.6 How next-token learning may suffer in our tasks
Much like in sophisticated creative tasks, in our tasks, the most natural way to generate the string is by planning various random latent choices (say $\bm{z}$) in advance and by producing a plan-conditioned distribution $p(\bm{s}|\bm{z})$ over coherent strings $\bm{s}$. However, next-token prediction (NTP) – or next-token learning to be precise – we argue, is myopic and may struggle to learn such a latent plan. Our argument extends that of B&N’24 to our even simpler tasks.
@@
Consider the sibling task where we must generate sibling-parent triplets $(\gamma,\gamma^{\prime},\Gamma)$. The most natural generative rule is to plan the last token (the parent) first and decide the children last. Think of this as learning a latent plan $\bm{z}:=\Gamma$. Then, learning the plan-conditioned generator $p(\gamma,\gamma^{\prime},\Gamma|\bm{z})$ factorizes to learning the distribution of children conditioned on a parent as $p(\gamma|\bm{z}:=\Gamma)$ and $p(\gamma^{\prime}|\bm{z}:=\Gamma)$ (due to conditional independence), and the trivial $p(\Gamma|\bm{z}:=\Gamma)$. This requires only as many parent-sibling edges as there are in the graph, i.e., $O(m\cdot n)$ many points, if there are $m$ parents, each with $n$ children. This is optimal.
@@
Things should proceed differently with NTP. We argue that a NTP-learner would fail to learn the plan $\bm{z}:=\Gamma$. The key intuition is that an NTP-learner learns the parent $\Gamma$ witnessing the siblings $(\gamma,\gamma^{\prime})$ as input. This is trivial to fit: the parent is simply the mutual neighbor of the two siblings revealed in the prefix; B&N’24 term this a “Clever Hans cheat” since the model witnesses and exploits part of the ground-truth it must generate (the siblings). Such cheats, being simpler than even the true generative rule, are quickly picked up during learning due to the well-known simplicity bias of gradient-descent-trained neural networks (Shah et al., 2020). Subsequently, any gradient supervision from the parent $\Gamma$ is lost (also known as gradient starvation (Pezeshki et al., 2021)), leaving no guidance to learn the latent plan, $z:=\Gamma$.
@@
After the Clever Hans cheat is learned, we conjecture that the NTP-learner learns the second sibling not through the plan-conditioned distribution $p(\gamma^{\prime}|\bm{z}:=\Gamma)$ but through the next-token-conditional, $p(\gamma^{\prime}|\gamma)$. This is a complex distribution: learning this may require witnessing every sibling-sibling pair totalling $O(m\cdot n^{2})$ many training data — larger by a factor of $n$ compared to the natural rule.
@@
Summary of argument. Abstractly, in a creative leap-of-thought task, it is most efficient to learn a well-planned random latent $p(\bm{z})$ and a subsequent latent-conditioned distribution $p(\bm{s}|\bm{z})$. However, NTP factorizes this into pieces of the form $p(s_{i}|\bm{s}_{<i},\bm{z})$. Consequently, on the later tokens, the model may be lured by Clever Hans cheats and learn uninformative latents. Conversely, the model may learn the earlier through complex rules, bereft of a latent plan. While this may not lead to complete breakdown of learning as in B&N’24 in our tasks, we hypothesize it must lead to data-hungry learning.